OpenStudy (anonymous):

Precalc question: I need help graphing the following equations. 1. x = 3 sin^3t y = 3 cos^3t 2. x = 7 sin t + sin 7t y = 7 cos t + cos 7t 3. x = 2t y = t + 5, -2 ≤ t ≤ 3 4. x = 2t – 1 y = t2 + 5, -4 ≤ t ≤ 4 5. x = 6 sin t y = 6 cos t, 0 ≤ t ≤ 2π

OpenStudy (anonymous):

Use a graphing utility. Wolframalpha will do. For number 1 http://www.wolframalpha.com/input/?i=parametricplot+%28+3+sin%28t%29%5E3+%2C3+cos%28t%29%5E3%29

OpenStudy (anonymous):

Mimic the example above for the rest

OpenStudy (anonymous):

The third one is line

OpenStudy (anonymous):

The 4th is a parabola and the 5th is a circle

OpenStudy (anonymous):

Are you with me?

OpenStudy (anonymous):

\[\frac{ x }{ 3 }=\sin ^3t,\sin t=\left( \frac{ x }{ 3 } \right)^{\frac{ 1 }{ 3 }}\] \[\cos t=\left( \frac{ y }{ 3 } \right)^{\frac{ 1 }{ 3 }}\] \[x ^{\frac{ 2 }{ 3 }}+y ^{\frac{ 2 }{ 3 }}=\sin ^2t+\cos ^2t=1\]